An aeroplane has a maximum acceleration of 6.0 m s–2. Calculate the minimum length for a straight runway so that the aeroplane, starting from rest, can safely take off at a velocity of 90 m s−1.


.



1
Expert's answer
2021-11-15T14:00:21-0500

"s=\frac{v^2-v_0^2}{2a}=\frac{90^2-0^2}{2\cdot 6}=675\ (m)" . Answer